OpenStudy (anonymous):

retailer is selling $120 mirror at 33% off. plus an additional 15% off what will the total cost be of the mirror?

OpenStudy (anonymous):

120*1/3=40 120-40=80 80*15/100 or 80*3/20=12 80-12=$68 The cost of the mirrow would be $68